OpenStudy (anonymous):

f '(pi/4) = cos(x) - xsin(x)

OpenStudy (anonymous):

wow. does this mean \[f'(x)=\cos(x)-x\sin(x)\] and you want \[f'(\frac{\pi}{4})\]?

OpenStudy (anonymous):

in other words you took the derivative of \[f(x)=x\cos(x)\] and now you want to evaluate at \[\frac{\pi}{4}\]

OpenStudy (anonymous):

yes that's what i tried to say

OpenStudy (anonymous):

yikes. asnwer is \[\frac{\sqrt{2}}{2}-\frac{\pi}{4}\frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

cos(pi/4)-(pi/4)sin(pi/4) (sqrt(2)/2)-(pi/4)(sqrt(2)/2)

OpenStudy (anonymous):

my computer takes forever to post an answer

OpenStudy (anonymous):

its okay but thank you very much for helping me